There are two jack screw assemblies both with an ISO Metric Trapezoidal ‘Tr 16 x 3’ screw, one single-thread and the other double-thread, to raise and lower a load of 450 kg. A ball-thrust bearing forms part of the assembly and therefore collar friction is negligible. The coefficient of sliding friction for both screws is 0.08.

(a) For each assembly, determine:

(i) the pitch, pitch diameter, lead and lead angle;

(ii) the running torque required to lower the load;

(iii) whether they are back-driveable or self-locking?

(b) Which assembly is the most efficient in lifting?

(c) One way of ensuring that a lifting device can be made to self-lock is to use a thrust collar rather than a ball-thrust bearing. For any of the assemblies found to be back-driveable in (a)(iii), what is the minimum collar diameter required to ensure it/they would self-lock?

(Assume the same value of friction coefficient of 0.08 in all situations).
Answer #1

the SCrews a Trapczeidel (a) Goth PitCk Nominal diamet Pitch du a mettia Nomined dra mete dp 6mm Lead = no. stads x Pitch, foKnow that torque require d to E (m co s α- s.no() 2. 4.57 cos φ = o-396 (same W: o·08 fox both) イ는 11 F- YSo g.BI = 441年.SNF= load ; 4414.SN cos φ, 0996 m44S (34 o.996 L = 10.467 L 26 o.99